OpenStudy (anonymous):

Use the divisibility rules to determine which of the following numbers 504 is divisible by. (Check all that apply.) 2 3 5 7 11

OpenStudy (danjs):

Decompose 504 into prime factors

OpenStudy (danjs):

2 * 252 ---to begin

OpenStudy (danjs):

2 * 2 * 2 * 63

OpenStudy (danjs):

2 * 2 * 2 * 3 * 21

OpenStudy (danjs):

2 * 2 * 2 * 3 * 3 * 7

OpenStudy (anonymous):

hm

OpenStudy (danjs):

2, 3, and 7 go into 504 evenly

OpenStudy (danjs):

or any combination of those prime factors

OpenStudy (danjs):

504 = 2 * 2 * 2 * 3 * 3 * 7

OpenStudy (danjs):

for example a combination, 2*2*3 = 12 will go into 504 even, or any combination of those prime factors

OpenStudy (danjs):

7*3 = 21 21 will evenly divide into 504

OpenStudy (danjs):

u get it?

OpenStudy (amistre64):

divisibility rules ... that like the sum of the digits is this or that type of rules?

OpenStudy (amistre64):

10 = -1 mod(11) 5(10)^2 +0(10)^1 +4(10)^0 5(-1)^2 +0(-1)^1 +4(-1)^0 mod(11) 5 +4 mod(11) 9 mod 11, since its not 0 mod 11 its not evenly divisible by 11

OpenStudy (amistre64):

10 = 0 mod 5 5(10)^2 +0(10)^1 +4(10)^0 5(0)^2 +0(0)^1 +4 mod(5) 4 mod(5), so unless it ends in a 0 or 5, then its not divisible by 5